# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

### v2

@@ 2.3. Disentangled Representation Learning
Disentangled representation learning, aiming to discern and separate fundamental explanatory factors (Bengio et al., 2013), stands as a pivotal method within deep learning. It strives to produce resilient, manageable, and interpretable representations. Existing research predominantly focuses on the field of CV, often employing methods like Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) to disentangle relevant latent factors (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021). Researchers explore techniques to align distinct representations with specific factors through explicit and implicit guidance, employing approaches like direct labeling, loss penalties, and regularization. For instance, DR-GAN (Tran et al., 2017) utilizes an adversarial network to generate diverse facial poses, aiming to disentangle pose information from facial attributes. In NLP, disentangled representation learning finds applications in various tasks like generation. For example, John et al. (John et al., 2018) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style. This separation allows precise control over content and style in text generation.
@@ 3.3.1. Citation-aware R-GCNs
For ”cites” and ”is cited by” edge subgraphs of paper nodes, inspired by Disruption-index (Wu et al., 2019), we introduce co-cited/citing strengths to integrate multi-hop citation information. The Disruption-index categorizes papers as disruptive or developmental based on the metric related to their betweenness centrality. Disruptive papers play a critical role in altering typical citation diffusion. We adapt this subgraph-focused metric to estimate node-pair closeness by considering co-cited/citing strengths. We aim to discern the varying significance of citation behaviors during message passing and distinguish crucial references and citations within ”is cited by” and ”cites” edges for the target paper. On one side, within ”is cited by” edges, when the source (reference) and target (citation) papers share more similar references, the influence of the source paper on the target paper might diminish. From the viewpoint of the target paper, the source paper could be considered as merely developmental, exerting a potentially negative impact on message passing. Moreover, since the information from this source paper is already conveyed through other original disruptive references, the negative co-citing effect might amplify attention toward these disruptive source papers. On the other side, within “cites” edges, if the source (citation) and target (reference) paper have more identical citations, they may share more commonness and are more closely connected, leading to a positive effect. Unlike co-citing relationships, citations are expansive and continuously growing, making it challenging to differentiate disruptive papers based on co-cited relationships alone. Therefore, we employ a normalized score of co-cited strength that considers both similarity and accumulated count to identify highly-related developmental citations. It’s crucial to note that the co-cited/citing strengths are dynamic scores that evolve over time. As new papers emerge, the values of normalized co-cited/citing strengths undergo changes not only due to the direct influence of new citations but also because of alterations in numerical values and maximum values. Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion. Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2021), to encode comparative and co-cited/citing features between paper nodes. To capture the discrepancy between source and target nodes, we encode the concatenation of their representations. Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network. Based on the adjacent matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes. Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016). For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$, we have:
@@
For other edges in the R-GCN, we apply GIN (Xu et al., 2018) with sum-pooling as the GNN encoder:
@@ 4.4. Evaluation Metric
[…]
@@
[…]
@@
[…]
@@ 5.1. Performance Comparison (RQ1)
[…]
@@
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
@@
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
@@
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
@@
[…]
@@ 5.2. Ablation Test (RQ2)
To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules. As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2021), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering. For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
@@ 5.3. Hyper-parameters Test (RQ3)
[…]

### v3

@@ 2.3. Disentangled Representation Learning
Disentangled representation learning, aiming to discern and separate fundamental explanatory factors (Bengio et al., 2013), stands as a pivotal method within deep learning. It strives to produce resilient, manageable, and interpretable representations. Existing research predominantly focuses on the field of CV, often employing methods like Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) to disentangle relevant latent factors (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021). Researchers explore techniques to align distinct representations with specific factors through explicit and implicit guidance, employing approaches like direct labeling, loss penalties, and regularization. For instance, DR-GAN (Tran et al., 2017) utilizes an adversarial network to generate diverse facial poses, aiming to disentangle pose information from facial attributes. In NLP, disentangled representation learning finds applications in various tasks like generation. For example, John et al. (John et al., 2019) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style. This separation allows precise control over content and style in text generation.
@@ 3.3.1. Citation-aware R-GCNs
For ”cites” and ”is cited by” edge subgraphs of paper nodes, inspired by Disruption-index (Wu et al., 2019), we introduce co-cited/citing strengths to integrate multi-hop citation information. The Disruption-index categorizes papers as disruptive or developmental based on the metric related to their betweenness centrality. Disruptive papers play a critical role in altering typical citation diffusion. We adapt this subgraph-focused metric to estimate node-pair closeness by considering co-cited/citing strengths. We aim to discern the varying significance of citation behaviors during message passing and distinguish crucial references and citations within ”is cited by” and ”cites” edges for the target paper. On one side, within ”is cited by” edges, when the source (reference) and target (citation) papers share more similar references, the influence of the source paper on the target paper might diminish. From the viewpoint of the target paper, the source paper could be considered as merely developmental, exerting a potentially negative impact on message passing. Moreover, since the information from this source paper is already conveyed through other original disruptive references, the negative co-citing effect might amplify attention toward these disruptive source papers. On the other side, within “cites” edges, if the source (citation) and target (reference) paper have more identical citations, they may share more commonness and are more closely connected, leading to a positive effect. Unlike co-citing relationships, citations are expansive and continuously growing, making it challenging to differentiate disruptive papers based on co-cited relationships alone. Therefore, we employ a normalized score of co-cited strength that considers both similarity and accumulated count to identify highly-related developmental citations. It’s crucial to note that the co-cited/citing strengths are dynamic scores that evolve over time. As new papers emerge, the values of normalized co-cited/citing strengths undergo changes not only due to the direct influence of new citations but also because of alterations in numerical values and maximum values. Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion. Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes. To capture the discrepancy between source and target nodes, we encode the concatenation of their representations. Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network. Based on the adjacent matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes. Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016). For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$, we have:
@@
For other edges in the R-GCN, we apply GIN (Xu et al., 2019) with sum-pooling as the GNN encoder:
@@ 4.4. Evaluation Metric
In our evaluation, we utilize two metrics: MALE (Mean Absolute Logarithmic Error) and LogR2 (Logarithmic R-squared). MALE calculates the absolute error between the target and predicted values post-logarithm transformation, reflecting the direct predictive capability:
@@
Regarding LogR2, it evaluates the coefficient of determination, calculated as the proportion of the dependent variable’s variation predictable from the independent variable, following logarithm transformation for both variables:
@@
LogR2 assumes a negative value when the model’s predictions fall short of the data’s mean value, indicating an inability to capture the underlying trend within the dataset. In our problem context, mirroring practical scenarios within dynamic contexts that require extrapolative abilities, encountering negative values is both frequent and expected due to inherent complexities. Consequently, by integrating MALE and LogR2, we holistically gauge the model’s accuracy in direct predictions for individual samples and its aptness in capturing broader trends across groups.
@@ 5.1. Performance Comparison (RQ1)
Table 2. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Computer Science. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
@@
#improve (%) 11.93 13.32 12.65 7.09 11.15 32.22 15.68 79.67
Table 3. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Chemistry. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
@@
#improve (%) 7.16 5.44 5.64 0.36 10.67 39.13 9.42 80.39
Table 4. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Psychology. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
@@
#improve (%) 14.29 13.96 13.86 8.87 11.92 44.38 10.38 134.40
(1) All models perform well on previous papers presented in the training set but experience degradation when dealing with fresh and immediate papers in the test set, particularly in terms of the LogR2. Many models fail to figure out the overall trend of immediate papers, which results in negative values in LogR2. Even the best baseline, MUCas, struggles with immediate papers due to limited graph structures, as it is a cascade prediction model designed for papers with sufficient citations. However, dynamic GNN models can effectively encode multi-hop information to complement the lack of cascade structures. Disentangled GNNs, in particular, effectively uncover crucial insights within both multi-hop graph structures and node representations, notably beneficial for datasets like Psychology, which are denser. Moreover, models such as SciBERT and HINTS leverage prior knowledge, like content information or citation distribution, to tackle these challenges. Comparatively, while SciBERT outperforms HINTS in fitting the increments of previous papers (SciBERT ¿ HINTS), it faces greater difficulty with fresh papers (SciBERT ¡ HINTS). This implies that graph structures exhibit better generality and transferability to unseen samples, while shared content information remains robust for immediate papers with limited graph structures, both alleviating issues related to data distribution shifts.
@@
(3) While Disentanglement methods offer performance improvements, their adaptation to dynamic heterogeneous graphs is essential to achieve optimal flexibility. Notably, DIDA exhibits superior outcomes than ROLAND in LogR2 for Chemistry, even when relying solely on homogeneous graphs. Moreover, within the densest dataset, Psychology, Disentangled GNN models showcase impressive performance. Among them, CAL outperforms both MUCas and ROLAND in predicting fresh and immediate papers. However, in other scenarios, apart from DIDA, these methods face challenges in competing with task-specific models and dynamic GNNs.
(4) Our proposed model surpasses all baselines, by effectively modeling the dynamic heterogeneous graph and incorporating task-adjusted disentanglement. This leads to an average improvement of about 10% in MALE across all categories, along with a substantial advancement of over 30% in LogR2 for fresh and immediate papers. First, compared with traditional citation/cascade prediction models, it exhibits superior generalization and practicability through the fusion of content and context information within citation networks, as well as the disentanglement power. Content information aids in discerning the impact of previous papers and fortifies robustness towards immediate papers. Moreover, its adeptness in handling fresh and immediate papers stems from the rich information embedded within graph structures. Furthermore, we consider the unique properties of citation networks to encode complex high-order structural and temporal information, thereby outperforming common dynamic models. The proposed CGE facilitates the identification of influential papers related to the target paper based on comparative and multi-hop information. More importantly, our model introduces interpretability from disentanglement models without compromising performance and further adapts to the task for enhanced improvement. It excels in predicting precise citation increments and approaches the genuine contribution of papers by disentangling values derived from information diffusion and collective conformity, thus providing insights into the papers’ authentic impact.
@@ 5.2. Ablation Test (RQ2)
To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules. As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2022), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering. For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
@@ 5.3. Hyper-parameters Test (RQ3)
Figure 4. Results of hyper-parameters test in terms of LogR2.